iosangel:

Directions: For each situation, you need to determine if the object has KE or GPE or both. Then find their values. Be sure to include the correct units. 1. A 2-kg bowling ball sits on top of a building that is 40 meters tall. Circle one: KE / GPE / both Show your work for finding the values of each type of energy the object has:

iosangel:

@Shadow

standfast01:

@Elsa213

standfast01:

Trust me shes good

Elsa213:

Not at physics, boi e.e

Elsa213:

GPE=mgh GPE= \(\bf 2 kg*9.8 m/s^2*40\) GPE= 19.6*40 GPE=784 joules(Energy is measured in joules) Since the bowling ball is sitting the kinetic energy is 0 Your answer is `GPE`.

Elsa213:

@iosangel

iosangel:

Thank you

Elsa213:

Ofc. c: <3